Each day, two out of the three teams in a class are randomly selected to participate in a MATHCOUNTS trial competition. What is the probability that Team A is selected on at least two of the next three days? Express your answer as a common fraction.
Answer by stanbon(75887)

Each day, two out of the three teams in a class are randomly selected to participate in a MATHCOUNTS trial competition. What is the probability that Team A is selected on at least two of the next three days? Express your answer as a common fraction.
--------------------
P(A is selected on a particular day) = 1-P(A is not selected)
= 1-{[2C2]/[3C2]}
= 1 - 1/3
= 2/3
----
Then, p(2<= x <=3) = 1- binomcdf(3,2/3,1) = 0.7407 = 20/27
==============================================================
cheers,
Stan H.